# Towards a theory of dissipative Dark Matter I: the Born limit

### v1

@@ 4.1 Emission factors and elastic scattering amplitudes
$\mathcal{E}_{{V}}=\sum_{n}\eta_{n}g_{n}\frac{p_{n}\cdot\epsilon^{*}(q)}{p_{n}\cdot q}.$ (4.1)
@@ 4.2 Soft dipole emission
Figure 5: Dipole $(n=1)$ and quadrupole $(n=2)$ energy differential cross-sections for distinguishable particle scattering with $v_{i}\approx 4\times 10^{-3}$. Solid lines are the massless limits of our analytical energy-differential cross section expressions; dotted lines show the associated soft emission limits. A separation scale between “hard” and “soft” emission at half the available CM energy $\Lambda\sim\mu v_{i}^{2}/4$ is introduced; soft modes ($\omega\lesssim\Lambda$) are shaded light purple, hard modes ($\omega\gtrsim\Lambda$) light blue. Left: short-range mediation. The brown line shows the dipole result (3.26). Quadrupole lines follow (3.36) (full) and (4.25) (soft). Right: long-range mediation. The brown line shows the dipole result (3.22), and the dotted grey line its soft limit (4.12b). For quadrupole emission, solid colored lines show the full results (3.30), and the dotted colored lines the soft limit (4.26).
@@ 4.3 Soft quadrupole emission: amplitudes
For long-range forces, we therefore introduce corrections detailed below, in order to retrieve an accurate $x\ll 1$ behavior of the soft energy differential cross-section. It results in quadrupole soft factors that differ between long-range and short-range mediation. Moreover, for both types of mediation, long- and short-range, the full kinematics [^10] A similar observation seems to be made in [104], where it was noted that the original Low theorem [100] is formulated in terms of the physical momenta of the emission process, without relying on the $|\mathbf{q}|\to 0$ limit of exchanged momenta. of the emission process must be retained in the derivation of the soft factors (4.1), (4.2) and (4.3), as finite $\mathbf{q}$-corrections become essential at quadrupole order.
@@
$T_{1}$ $\equiv|\vec{v}_{f}|^{4}+|\vec{v}_{i}|^{4}+|\vec{v}_{i}|^{2}|\vec{v}_{f}|^{2}\left(2-4(\vec{v}_{i}\cdot\vec{v}_{f})^{2}\right),$ (4.14)
@@
To derive the soft quadrupole emission amplitude, we proceed by taking the quadrupole soft factor (4.19) and multiplying it by the elastic cross-sections (4.5) (distinguishable particle scattering) and (4.7) (identical particle scattering). In both cases, the resulting soft expression reproduces the full quadrupole emission amplitudes, namely,
@@
Before presenting the soft energy differential cross-sections associated with the squared amplitudes derived in this section, we examine their counterparts in the case of long-range mediated interactions.
@@
which is in agreement with [63]. For distinguishable particle scattering, multiplying the soft factor (4.22) by the Rutherford squared amplitude (4.4) leads, as in the short-range mediation case, to the equality (4.16) between the soft emission amplitude and the full result (3.5). In the case of identical particle scattering, the soft emission amplitude is readily obtained by multiplying (4.22) with the elastic matrix element squared (4.6). Although this product correctly captures the $x\ll 1$ behavior of the full emission amplitudes (F.33), it cannot reproduce the full expressions, specifically the interference terms between the $t$- and $u$-channel.
@@
In contrast to short-rage mediation, it is crucial to carry along the $\mathcal{O}(q^{0})$-term in (4.3) in order to obtain the correct $x\ll 1$ behavior of the full NR amplitude; see App. D for a detailed discussion.
With these factors at hand, we now turn obtain the soft quadrupole emission amplitudes. In the case of distinguishable particle scattering, each amplitude is obtained by multiplying the quadrupole soft factors (4.23) and (4.24) with the squared Rutherford matrix element (4.4). These two products reproduce, respectively, the quadrupole emission amplitude (C.5) and the full result (C.2) in the limit $\lambda_{j}\ll(A_{j}/m_{j})^{2}$, and the full quadrupole amplitude (3.5), thereby yielding the equalities (4.18) and (4.20). For identical particle scattering, the situation parallels the other scenarios. Multiplying the quadrupole soft factor (4.23) with the squared Coulomb amplitude (4.6) reproduces the full quadrupole amplitude (F.35) (for $\kappa=0$), while using the soft factor (4.24) instead reproduces (F.34). In both cases, the interference between the $t$- and $u$-channel contributions is not fully captured, although the results coincide with the full expressions in the $x\ll 1$ limit.
@@
4.4 Soft quadrupole emission: energy differential cross-sections
We are now in a position to obtain the energy-differential cross-sections for quadrupole emission. For identical particle scattering, these are the leading order in $v_{i}$ and $r$ results. As mentioned several times, for non-identical particles, they are sub-leading to the dipole order, unless the coupling-to-mass ratios in the dipole factors $\mathcal{F_{D}}$ in (3.16) for long-range mediation or $\mathcal{F_{D}^{\prime}}$ in (3.24)-(3.25) for short-range mediation, happen to cancel.
@@ 4.4 Soft quadrupole emission: energy differential cross-sections
In the case of distinguishable particle scattering, we perform the phase-space integrations in (3.13) in the $x\ll 1$ limit, using the soft quadrupole emission amplitudes defined in Sec. 4.3. At leading order in $x$, this yields the following soft energy-differential cross-sections for short-range mediation,
@@
where the parameters $\alpha$ and the pre-factors $\mathcal{F}^{\prime}_{\mathcal{Q}}$ are defined in (3.33)-(3.35). While, in the case of long-range mediation, the corresponding soft energy differential cross-sections read,
@@
We note that both (4.25) and (4.26) agree at leading order in $x$ with the full quadrupole energy differential cross-sections, respectively (3.36) and (3.30), for all the scenarios considered in this work. In Fig. 5, as for the dipole emission case, we compare the soft and the full expressions, both for short-range (left panel) and long-range (right panel) mediation.
@@
These soft energy differential cross-sections agree with their full quadrupole counterparts, (3.38) and (3.41), for $\kappa=0$, once the latter are expanded to first order in $x$.
Figure 6: Energy differential cross-section for identical particle scattering with $v_{i}\approx 4\times 10^{-3}$ and $a$ corresponding to the couplings $g,y$ or $A/m$ depending on the dSIDM scenario considered. A separation scale between “hard” and “soft” emission at half the available CM energy $\Lambda\sim\mu v_{i}^{2}/4$ is introduced; soft modes ($\omega\lesssim\Lambda$) are shaded light purple, hard modes ($\omega\gtrsim\Lambda$) light blue. The solid colored lines represent the massless emission limit ($\kappa=0$) of the energy differential cross-sections from Sec. 3.5.2, while the colored dotted lines correspond to (4.27) (vector emission) and to (4.28)-(4.29) (scalar emission), obtained using the soft quadrupole emission amplitude squared corrected by the factor (4.21). Light grey dotted lines show the uncorrected soft approximation. Left: vector emission. Right: scalar emission.
@@
which agrees with the full quadrupole counterpart (3.43), for $\kappa=0$, when expanded to leading order in $x$. Similarly, for fermionic DM candidates, we obtain the soft energy differential cross-section
@@
Finally, in the case of short-range mediation, the resulting soft energy differential cross-sections are in perfect agreement with the full results ($\kappa=0$) presented in Sec. 3.5.2, once the latter are expanded at first order in $x$. Since the resulting expressions can be directly inferred from the full results, we refrain from presenting them explicitly.
@@ 5 Energy loss rate
Using a standard reduction of the phase space integrals [^11] This reduction holds under the assumption that the CM frame and the medium frame coincide in the NR framework [105]. It introduces an error of order $\mathcal{O}(\mathbf{v}_{CM})$, with $\mathbf{v}_{CM}$ the velocity of the CM relative to the medium frame., the energy loss rate (5.1) can then be brought to an integral over the CM energy-differential cross-sections derived above,
@@
The energy loss rate can be evaluated analytically for massless emission using (5.3). For the emission of a massive particle, however, a numerical evaluation is required. In this case, for dipole emission and the scattering of identical particles, the energy differential cross-section, defined in Sec. 3.4 and Sec. 3.5.2, must be inserted in (5.3), while for quadrupole emission from the scattering of distinguishable particles, the energy differential cross-section provided in App. F must be used.
@@ 5.1 Scattering of distinguishable particles
In the case of interactions mediated by a long-range force, by substituting the energy differential cross-sections (3.22) (for $m_{\phi,V}=0$) and (3.30) in (5.3), we obtain for the fermionic dSIDM options, $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}\phi/V$, $\tilde{\chi}_{1}\tilde{\chi}_{2}\to\tilde{\chi}_{1}\tilde{\chi}_{2}\phi$, as well as for the scalar dSIDM option $S_{1}S_{2}\to S_{1}S_{2}\phi/V$, and $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$ under the premise $(A_{j}/m_{j})^{2}\gg\lambda_{j}$, the emission rate
@@
We now evaluate the energy loss rate arising from short-range mediated forces, using the energy-differential cross sections given in (3.26) for dipole and (3.36) for quadrupole emission. For the various dSIDM scenarios considered, including the processes $S_{1}S_{2}\to S_{1}S_{2}\phi$ and $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$ in the limit $\langle|\mathcal{M}_{\text{vertex}}|^{2}\rangle\ll\langle|\mathcal{M}_{\text{external leg}}|^{2}\rangle$ (see App. C), we obtain
@@ 5.1.1 Dipole emission
If the emitted particle is massive, the energy loss rate (5.1) must be evaluated numerically using the dipole energy differential cross-sections (3.22) (long-range mediation) and (3.26) (short-range mediation). As discussed in Sec. 3.4, we emphasize that in the case of dipole emission, the form of the $m_{\phi,V}$-dependence appearing in the unpolarized emission amplitudes is the same for both types of mediation and depends only on the nature of the emitted particle, not on the spin of the DM candidates.
@@
In the upper panels of Fig. 7, we plot the ratio $R$ defined in (5.5) which compares the energy-loss rates between massless and massive emissions. For both types of mediation, the difference in the effects of the emitted particle mass between $\phi$- and $V$-emission is mild up to $m_{\phi,V}\sim T$. When $m_{\phi,V}=3T$, the mass of the emitted particle already exceeds the typical kinetic energy of the plasma, and the energy-loss rate is exponentially suppressed. In this regime, the energy loss rate from massive $\phi$-emission is more suppressed than in the case of massive $V$-emission. Additionally, we note that the impact of $m_{\phi,V}$ on the energy loss rate of short-range mediation is weaker by an order of magnitude compared to the long-range mediation scenario.
@@ 5.1.2 Quadrupole emission
Higher powers in $T$ of the energy loss rates, $T^{3/2}$ for long-range mediation in (5.6) and $T^{7/2}$ for short-range mediation in (5.8), arise from quadrupole emission. Their suppression relative to dipole emission scales with the typical kinetic energy in the system, $T\sim\mu\langle v_{i}^{2}\rangle$. At this order, the energy loss rates become model-dependent for both mediations, reflecting the structure of the corresponding energy differential cross-sections (3.30) and (3.36). This model dependence can be qualitatively evaluated by comparing the energy loss rates across different dSIDM scenarios.
@@
Starting with long-range mediated processes, $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}V$, $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}\phi$, $\tilde{\chi}_{1}\tilde{\chi}_{2}\to\tilde{\chi}_{1}\tilde{\chi}_{2}\phi$, and $S_{1}S_{2}\to S_{1}S_{2}V$ all give rise to quadrupole radiation. In the first three cases, the dominant contribution arises from external leg emission, while in the last process, there is an additional vertex emission contribution. For the processes $S_{1}S_{2}\to S_{1}S_{2}\phi$ and $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$, two limits must be considered. In the limit $(A_{j}/m_{j})^{2}\gg\lambda_{j}$ the latter processes are dominated by external leg emission, yielding quadrupole emission as well. Consequently, by assuming a common reduced mass across all the DM candidates and massless emission, the ratios of the energy loss rates between the different dSIDM scenarios reduce to
@@
Here, the subscript $\mathcal{Q}$ indicates the quadrupole nature of the emission. The index $l,k$ label the scenarios, with the shorthand notation listing the emitted particle first, followed by the DM candidates; for the latter, one may, individually for $l$ and/or $k$, chose either the unbracketed or bracketed DM candidate for the equation to hold. When the processes $S_{1}S_{2}\to S_{1}S_{2}\phi$ and $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$ are dominated by vertex emission, i.e., in the regime $(A_{j}/m_{j})^{2}\ll\lambda_{j}$, the energy loss rate ratio becomes
@@
with $C_{k}\approx 42.7$ for $k=\phi\,\chi_{1,2}(\tilde{\chi}_{1,2})$, and $C_{k}\approx 53.4$ for $k=V\,\chi_{1,2}({S}_{1,2})$.
@@
For massive emission, we evaluate numerically the energy loss rate using the energy differential cross-sections defined in App. F. We plot the corresponding ratio $R$, as defined in (5.5), in the lower panels of Fig. 7.
@@
  • For dipole emission, we plot the rate corresponding to $\phi$-emission for both types of mediation. Although the difference with $V$-emission remains small, this emission type presents the most efficient suppression when the emitted particle is massive.
  • For quadrupole emission, we have shown that model-dependent differences remain mild in both the long-range and short-range cases for massless emission; see (5.11) and (5.13), respectively. For massive emission, the same conclusion holds in the case of long-range mediation. In contrast, for short-range mediation, small differences start to arise for $m_{\phi,V}\sim T$, with the largest suppression observed for $V$-emission. We therefore use the latter as a representative example.
Finally, we emphasize that adding a small but finite mass to the emitted particle provides a natural way to modulate the strength of the dissipative effects across different halo sizes. Once typical CM energy—associated with the typical mean spead of DM particles—lies below the emitted particle mass scale, the dissipative effects are strongly suppressed.

### v2

@@ 4.1 Emission factors and elastic scattering amplitudes
$\mathcal{E}_{{V}}=\sum_{n}\eta_{n}g_{n}\frac{p_{n}\cdot\varepsilon^{*}(q)}{p_{n}\cdot q}.$ (4.1)
@@ 4.2 Soft dipole emission
Figure 5: Dipole $(n=1)$ and quadrupole $(n=2)$ energy-differential cross sections for distinguishable particle scattering. Solid lines are the massless limits of our analytical energy-differential cross section expressions; dotted lines show the associated soft emission limits. A separation scale between “hard” and “soft” emission at half the available CM energy $\Lambda\sim\mu v_{i}^{2}/4$ is introduced; soft modes ($\omega\lesssim\Lambda$) are shaded light purple, hard modes ($\omega\gtrsim\Lambda$) light blue. Left: short-range mediation. The brown line shows the dipole result (3.32), while the brown dotted line shows its soft limit (4.12a). Quadrupole lines follow (3.41) (full) and (4.25) (soft). Right: long-range mediation. The brown line shows the dipole result (3.28), and the corresponding dotted line shows its soft limit (4.12b). For quadrupole emission, solid colored lines show the full results (3.35), and the dotted colored lines the soft limit (4.26).
@@ 4.3 Soft quadrupole emission: amplitudes
For long-range forces, we therefore introduce corrections detailed below, in order to retrieve an accurate $x\ll 1$ behavior of the soft energy-differential cross section. It results in quadrupole soft factors that differ between long-range and short-range mediation. Moreover, for both types of mediation, long- and short-range, the full kinematics [^11] A similar observation seems to be made in [111], where it was noted that the original Low theorem [107] is formulated in terms of the physical momenta of the emission process, without relying on the $|\mathbf{q}|\to 0$ limit of exchanged momenta. of the emission process must be retained in the derivation of the soft factors (4.1), (4.2) and (4.3), as finite $\mathbf{q}$-corrections become essential at quadrupole order.
@@
$T_{1}$ $\equiv|\vec{v}_{f}|^{4}+|\vec{v}_{i}|^{4}+|\vec{v}_{i}|^{2}|\vec{v}_{f}|^{2}\left(2-4(\hat{\vec{v}}_{i}\cdot\hat{\vec{v}}_{f})^{2}\right),$ (4.14)
@@
To derive the soft quadrupole emission amplitude, we proceed by taking the quadrupole soft factor (4.19) and multiplying it by the elastic amplitudes squared (4.5) (distinguishable particle scattering) and (4.7) (identical particle scattering). In both cases, the resulting soft expression reproduces the full quadrupole emission amplitudes, namely,
@@
Before presenting the soft energy-differential cross sections associated with the squared amplitudes derived in this section, we examine their counterparts in the case of long-range mediated interactions.
@@
which is in agreement with [75]. For distinguishable particle scattering, multiplying the soft factor (4.22) by the Rutherford squared amplitude (4.4) leads, as in the short-range mediation case, to the equality (4.16) between the soft emission amplitude and the full result (3.5.1). In the case of identical particle scattering, the soft emission amplitude is readily obtained by multiplying (4.22) with the elastic matrix element squared (4.6). Although this product correctly captures the $x\ll 1$ behavior of the full emission amplitudes (F.33), it cannot reproduce the full expressions, specifically the interference terms between the $t$- and $u$-channel.
@@
In contrast to short-range mediation, it is crucial to carry along the $\mathcal{O}(q^{0})$-term in (4.3) in order to obtain the correct $x\ll 1$ behavior of the full NR amplitude; see App. D for a detailed discussion.
With these factors at hand, we now obtain the soft quadrupole emission amplitudes. In the case of distinguishable particle scattering, each amplitude is obtained by multiplying the quadrupole soft factors (4.23) and (4.24) with the squared Rutherford matrix element (4.4). These two products reproduce, respectively, the quadrupole emission amplitude (C.5) and the full result (C.2) in the limit $\lambda_{j}\ll(A_{j}/m_{j})^{2}$, and the full quadrupole amplitude (3.5.1), thereby yielding the equalities (4.18) and (4.20). For identical particle scattering, the situation parallels the other scenarios. Multiplying the quadrupole soft factor (4.23) with the squared Coulomb amplitude (4.6) reproduces the full quadrupole amplitude (F.35) (for $\kappa=0$), while using the soft factor (4.24) instead reproduces (F.34), up to the interferences between the $t$- and $u$-channel contributions, which are not fully captured. Nevertheless, the results coincide with the full expressions in the $x\ll 1$ limit.
@@
4.4 Soft quadrupole emission: energy-differential cross sections
We are now in a position to obtain the energy-differential cross sections for quadrupole emission. For identical particle scattering, these are the leading order in $v_{i}$ and $r$ results. As mentioned several times, for non-identical particles, they are sub-leading to the dipole order, unless the coupling-to-mass ratios in the dipole factors $\mathcal{F_{D}}$ in (3.22) for long-range mediation or $\mathcal{F_{D}^{\prime}}$ in (3.30)-(3.31) for short-range mediation, happen to cancel.
@@ 4.4 Soft quadrupole emission: energy-differential cross sections
In the case of distinguishable particle scattering, we perform the phase-space integrations in (3.19) in the $x\ll 1$ limit, using the soft quadrupole emission amplitudes defined in Sec. 4.3. At leading order in $x$, this yields the following soft energy-differential cross sections for short-range mediation,
@@
where the parameters $\alpha$ and the pre-factors $\mathcal{F}^{\prime}_{\mathcal{Q}}$ are defined in (3.38)-(3.40). While, in the case of long-range mediation, the corresponding soft energy-differential cross sections read,
@@
We note that both (4.25) and (4.26) agree at leading order in $x$ with the full quadrupole energy-differential cross sections, respectively (3.41) and (3.35), for all the scenarios considered in this work. In Fig. 5, as for the dipole emission case, we compare the soft and the full expressions, both for short-range (left panel) and long-range (right panel) mediation.
@@
These soft energy-differential cross sections agree with their full quadrupole counterparts, (3.43) and (3.46), for $\kappa=0$, once the latter are expanded to first order in $x$.
Figure 6: Energy-differential cross section for identical particle scattering mediated by a long-range interaction. Here, $a$ corresponds to the couplings $g,y$ or $A/m$ depending on the dSIDM scenario considered. A separation scale between “hard” and “soft” emission at half the available CM energy $\Lambda\sim\mu v_{i}^{2}/4$ is introduced; soft modes ($\omega\lesssim\Lambda$) are shaded light purple, hard modes ($\omega\gtrsim\Lambda$) light blue. The solid colored lines represent the massless emission limit ($\kappa=0$) of the energy-differential cross sections from Sec. 3.5.2, while the colored dotted lines correspond to (4.27) (vector emission) and to (4.28)-(4.29) (scalar emission), obtained using the soft quadrupole emission amplitude squared corrected by the factor (4.21). The light grey dotted lines show the uncorrected soft approximation. Left: vector emission. Right: scalar emission.
@@
which agrees with the full quadrupole counterpart (3.48), for $\kappa=0$, when expanded to leading order in $x$. Similarly, for fermionic DM candidates, we obtain the soft energy-differential cross section
@@
Finally, in the case of short-range mediation, the resulting soft energy-differential cross sections are in perfect agreement with the full results ($\kappa=0$) presented in Sec. 3.5.2, once the latter are expanded at first order in $x$. Since the resulting expressions can be directly inferred from the full results, we refrain from presenting them explicitly.
@@ 5 Energy loss rate
Using a standard reduction of the phase space integrals [^12] This reduction holds under the assumption that the CM frame and the medium frame coincide in the NR framework [112]. It introduces an error of order $\mathcal{O}(\mathbf{v}_{\text{CM}})$, with $\mathbf{v}_{\text{CM}}$ the velocity of the CM relative to the medium frame., the energy loss rate (5.1) can then be brought to an integral over the CM energy-differential cross sections derived above,
@@
The energy loss rate can be evaluated analytically for massless emission using (5.3). For the emission of a massive particle, however, a numerical evaluation is required. In this case, for dipole emission and the scattering of identical particles, the energy-differential cross section, defined in Sec. 3.4 and Sec. 3.5.2, must be inserted in (5.3), while for quadrupole emission from the scattering of distinguishable particles, the energy-differential cross section provided in App. F must be used.
@@ 5.1 Scattering of distinguishable particles
In the case of interactions mediated by a long-range force, by substituting the energy-differential cross sections (3.28) (for $m_{\phi,V}=0$) and (3.35) in (5.3), we obtain for the fermionic dSIDM options, $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}\phi/V$, $\tilde{\chi}_{1}\tilde{\chi}_{2}\to\tilde{\chi}_{1}\tilde{\chi}_{2}\phi$, as well as for the scalar dSIDM option $S_{1}S_{2}\to S_{1}S_{2}\phi/V$, and $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$ under the premise $(A_{j}/m_{j})^{2}\gg\lambda_{j}$, the emission rate
@@
We now evaluate the energy loss rate arising from short-range mediated forces, using the energy-differential cross sections given in (3.32) (for $m_{\phi,V}=0$) for dipole and (3.41) for quadrupole emission. For the various dSIDM scenarios considered, including the processes $S_{1}S_{2}\to S_{1}S_{2}\phi$ and $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$ in the limit $\langle|\mathcal{M}_{\text{vertex}}|^{2}\rangle\ll\langle|\mathcal{M}_{\text{external leg}}|^{2}\rangle$ (see App. C), we obtain
@@ 5.1.1 Dipole emission
If the emitted particle is massive, the energy loss rate (5.3) must be evaluated numerically using the dipole energy-differential cross sections (3.28) (long-range mediation) and (3.32) (short-range mediation). As discussed in Sec. 3.4, we emphasize that in the case of dipole emission, the form of the $m_{\phi,V}$-dependence appearing in the unpolarized emission amplitudes is the same for both types of mediation and depends only on the nature of the emitted particle, not on the spin of the DM candidates.
@@
In the upper panels of Fig. 7, we plot the ratio $R$ defined in (5.5) which compares the energy loss rates between massless and massive emissions. For both types of mediation, the difference in the effects of the emitted particle mass between $\phi$- and $V$-emission is mild up to $m_{\phi,V}\sim T$. When $m_{\phi,V}=3T$, the mass of the emitted particle already exceeds the typical kinetic energy of the plasma, and the energy loss rate is exponentially suppressed. In this regime, the energy loss rate from massive $\phi$-emission is more suppressed than in the case of massive $V$-emission. Additionally, we note that the impact of $m_{\phi,V}$ on the energy loss rate of short-range mediation is weaker by an order of magnitude compared to the long-range mediation scenario.
@@ 5.1.2 Quadrupole emission
Higher powers in $T$ of the energy loss rates, $T^{3/2}$ for long-range mediation in (5.6) and $T^{7/2}$ for short-range mediation in (5.8), arise from quadrupole emission. Their suppression relative to dipole emission scales with the typical kinetic energy in the system, $T\sim\mu\langle v_{i}^{2}\rangle$. At this order, the energy loss rates become model-dependent for both mediations, reflecting the structure of the corresponding energy-differential cross sections (3.35) and (3.41). This model dependence can be qualitatively evaluated by comparing the energy loss rates across different dSIDM scenarios.
@@
Starting with long-range mediated processes, $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}V$, $\chi_{1}\chi_{2}\to\chi_{1}\chi_{2}\phi$, $\tilde{\chi}_{1}\tilde{\chi}_{2}\to\tilde{\chi}_{1}\tilde{\chi}_{2}\phi$, and $S_{1}S_{2}\to S_{1}S_{2}V$ all give rise to quadrupole radiation. In the first three cases, the dominant contribution arises from external leg emission, while in the last process, there is an additional vertex emission contribution. For the processes $S_{1}S_{2}\to S_{1}S_{2}\phi$ and $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$, two limits must be considered. In the limit $(A_{j}/m_{j})^{2}\gg\lambda_{j}$, the latter processes are dominated by external leg emission, yielding quadrupole emission as well. Consequently, by assuming a common reduced mass across all the DM candidates and massless emission, the ratios of the energy loss rates between the different dSIDM scenarios reduce to
@@
Here, the subscript $\mathcal{Q}$ indicates the quadrupole nature of the emission. The indices $l,k$ label the scenarios, with the shorthand notation listing the emitted particle first, followed by the DM candidates; for the latter, one may, individually for $l$ and/or $k$, choose either the unbracketed or bracketed DM candidate for the equation to hold. When the processes $S_{1}S_{2}\to S_{1}S_{2}\phi$ and $\tilde{S}_{1}\tilde{S}_{2}\to\tilde{S}_{1}\tilde{S}_{2}\phi$ are dominated by vertex emission, i.e., in the regime $(A_{j}/m_{j})^{2}\ll\lambda_{j}$, the energy loss rate ratio becomes
@@
with $C_{k}=128/3$ for $k=\phi\,\chi_{1,2}(\tilde{\chi}_{1,2})$, and $C_{k}=160/3$ for $k=V\,\chi_{1,2}({S}_{1,2})$.
@@
For massive emission, we evaluate numerically the energy loss rate using the energy-differential cross sections defined in App. F. We plot the corresponding ratio $R$, as defined in (5.5), in the lower panels of Fig. 7.
@@
  • For dipole emission, we plot the rate corresponding to $\phi$-emission for both types of mediation, as it undergoes the strongest suppression when the emitted particle is massive, despite the small difference with $V$-emission.
  • For quadrupole emission, we have shown that model-dependent differences remain mild in both the long-range and short-range cases for massless emission; see (5.11) and (5.13), respectively. For massive emission, the same conclusion holds for long-range mediation. In contrast, for short-range mediation, small differences start to arise for $m_{\phi,V}\sim T$, with the largest suppression observed for $V$-emission. We therefore use the latter as a representative example.
Finally, we emphasize that adding a small but finite mass to the emitted particle provides a natural way to modulate the strength of the dissipative effects across different halo sizes. Once typical CM energy—associated with the typical mean speed of DM particles—lies below the emitted particle mass scale, the dissipative effects are strongly suppressed.